OpenStudy (daggiemib):

A blimp provides aerial television views of a football game. The television camera sights the stadium at a 7degrees angleof depression. The altitude of the blim is 400 meters. Which of the following is the line-of-sight distance from the television camera to the base of the stadium? Round to the nearest meter. 403m 3,257m 3,282m 4,500m

OpenStudy (nuttyliaczar):

Using what we know, you can make a right triangle and solve for the hypotenuse

OpenStudy (daggiemib):

cos(7)=400/x

OpenStudy (daggiemib):

the calculator said no solution

OpenStudy (nuttyliaczar):



OpenStudy (nuttyliaczar):

You should use the 83 degrees instead, since the 7 is just there to tell us what the angle for the relevant triangle is

OpenStudy (nuttyliaczar):

\[\cos \theta=adjacent/hypotenuse\]

OpenStudy (daggiemib):

all my calculators arent working

OpenStudy (nuttyliaczar):

You get 3282m when you do \[hypotenuse=adjacent/\cos \theta\]

OpenStudy (anonymous):

try evaluating the cosine of the angle and then solving

OpenStudy (nuttyliaczar):

I figured that giving the answer after knowing that much wasn't too much

OpenStudy (daggiemib):

no i get it thank you

OpenStudy (nuttyliaczar):

And if you're ever confused about which angle, think about what values cosine takes with different angles

OpenStudy (nuttyliaczar):

If you have a really slim angle of depression like here, your cosine value will be really tiny

OpenStudy (nuttyliaczar):

And likewise, if your angle of depression was almost 90, it would be close to 1